हिंदी

Two Unbiased Dice Are Thrown. Find the Probability That: (I) Neither a Doublet Nor a Total of 8 Will Appear

Advertisements
Advertisements

प्रश्न

Two unbiased dice are thrown. Find the probability that  neither a doublet nor a total of 8 will appear

Advertisements

उत्तर

We know that in a single throw of two dices, the total number of possible outcomes is (6 × 6) = 36.
Let S be the sample space.
Then n(S) = 36

Let E1 = event of getting neither a doublet nor a total of 8
Then E1' = event of getting either a doublet or a total of 8
∴ E1' = {(1, 1), (2, 2), (3, 3), (4, 4), (5, 5), (6, 6), (2, 6), (3, 5), (5, 3), (6, 2)}
i.e. n(E1') = 10
Thus, P(E1') = \[\frac{10}{36}\]

Hence, required probability P(E1) = 1- P(E1')

\[= 1 - \frac{10}{36} = \frac{36 - 10}{36} = \frac{26}{36} = \frac{13}{18}\]
 

 

shaalaa.com
  क्या इस प्रश्न या उत्तर में कोई त्रुटि है?
अध्याय 33: Probability - Exercise 33.3 [पृष्ठ ४७]

APPEARS IN

आर.डी. शर्मा Mathematics [English] Class 11
अध्याय 33 Probability
Exercise 33.3 | Q 28.1 | पृष्ठ ४७

वीडियो ट्यूटोरियलVIEW ALL [1]

संबंधित प्रश्न

A card is selected from a pack of 52 cards.

  1. How many points are there in the sample space?
  2. Calculate the probability that the card is an ace of spades.
  3. Calculate the probability that the card is
    1. an ace
    2. black card.

There are four men and six women on the city council. If one council member is selected for a committee at random, how likely is it that it is a woman?


A fair coin is tossed four times, and a person win Re 1 for each head and lose Rs 1.50 for each tail that turns up.

From the sample space calculate how many different amounts of money you can have after four tosses and the probability of having each of these amounts.


Three coins are tossed once. Find the probability of getting

  1. 3 heads
  2. 2 heads
  3. at least 2 heads
  4. at most 2 heads
  5. no head
  6. 3 tails
  7. exactly two tails
  8. no tail
  9. atmost two tails.

Check whether the following probabilities P(A) and P(B) are consistently defined

P(A) = 0.5, P(B) = 0.7, P(A ∩ B) = 0.6


Check whether the following probabilities P(A) and P(B) are consistently defined

P(A) = 0.5, P(B) = 0.4, P(A ∪ B) = 0.8


Fill in the blank in following table:

P(A) P(B) P(A ∩ B) P(A ∪ B)
0.5 0.35 .... 0.7

A and B throw a pair of dice. If A throws 9, find B's chance of throwing a higher number.

 

Two unbiased dice are thrown. Find the probability that the sum of the numbers obtained on the two dice is neither a multiple of 2 nor a multiple of 3


A bag contains 8 red, 3 white and 9 blue balls. If three balls are drawn at random, determine the probability that  all the balls are of different colours.


Which of the cannot be valid assignment of probability for elementary events or outcomes of sample space S =  {w1, w2, w3, w4, w5, w6, w7}:

Elementary events: w1 w2 w3 w4 w5 w6 w7
(i) 0.1 0.01 0.05 0.03 0.01 0.2 0.6

Which of the cannot be valid assignment of probability for elementary events or outcomes of sample space S =  {w1, w2, w3, w4, w5, w6, w7}:

Elementary events: w1 w2 w3 w4 w5 w6 w7
(ii)
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]

Which of the cannot be valid assignment of probability for elementary events or outcomes of sample space S =  {w1, w2, w3, w4, w5, w6, w7}:

Elementary events: w1 w2 w3 w4 w5 w6 w7
(iii) 0.7 0.06 0.05 0.04 0.03 0.2 0.1

Which of the cannot be valid assignment of probability for elementary events or outcomes of sample space S =  {w1, w2, w3, w4, w5, w6, w7}:

Elementary events: w1 w2 w3 w4 w5 w6 w7
(iv)
\[\frac{1}{14}\]
\[\frac{2}{14}\]
\[\frac{3}{14}\]
\[\frac{4}{14}\]
\[\frac{5}{14}\]
\[\frac{6}{14}\]
\[\frac{15}{14}\]

An urn contains twenty white slips of paper numbered from 1 through 20, ten red slips of paper numbered from 1 through 10, forty yellow slips of paper numbered from 1 through 40, and ten blue slips of paper numbered from 1 through 10. If these 80 slips of paper are thoroughly shuffled so that each slip has the same probability of being drawn. Find the probabilities of drawing a slip of paper that is blue or white


An urn contains twenty white slips of paper numbered from 1 through 20, ten red slips of paper numbered from 1 through 10, forty yellow slips of paper numbered from 1 through 40, and ten blue slips of paper numbered from 1 through 10. If these 80 slips of paper are thoroughly shuffled so that each slip has the same probability of being drawn. Find the probabilities of drawing a slip of paper that is red or yellow and numbered 1, 2, 3 or 4


An urn contains twenty white slips of paper numbered from 1 through 20, ten red slips of paper numbered from 1 through 10, forty yellow slips of paper numbered from 1 through 40, and ten blue slips of paper numbered from 1 through 10. If these 80 slips of paper are thoroughly shuffled so that each slip has the same probability of being drawn. Find the probabilities of drawing a slip of paper that is numbered 5, 15, 25, or 35


In a leap year the probability of having 53 Sundays or 53 Mondays is ______.


Three-digit numbers are formed using the digits 0, 2, 4, 6, 8. A number is chosen at random out of these numbers. What is the probability that this number has the same digits?


One mapping (function) is selected at random from all the mappings of the set A = {1, 2, 3, ..., n} into itself. The probability that the mapping selected is one to one is ______.


If the letters of the word ALGORITHM are arranged at random in a row what is the probability the letters GOR must remain together as a unit?


Six new employees, two of whom are married to each other, are to be assigned six desks that are lined up in a row. If the assignment of employees to desks is made randomly, what is the probability that the married couple will have nonadjacent desks?


Four candidates A, B, C, D have applied for the assignment to coach a school cricket team. If A is twice as likely to be selected as B, and B and C are given about the same chance of being selected, while C is twice as likely to be selected as D, what are the probabilities that C will be selected?


The accompanying Venn diagram shows three events, A, B, and C, and also the probabilities of the various intersections (for instance, P(A ∩ B) = .07). Determine P(A ∪ B)


A bag contains 8 red and 5 white balls. Three balls are drawn at random. Find the probability that all the three balls are red


If the letters of the word ASSASSINATION are arranged at random. Find the probability that four S’s come consecutively in the word


If the letters of the word ASSASSINATION are arranged at random. Find the probability that two I’s and two N’s come together


If the letters of the word ASSASSINATION are arranged at random. Find the probability that no two A’s are coming together


In a non-leap year, the probability of having 53 tuesdays or 53 wednesdays is ______.


Three numbers are chosen from 1 to 20. Find the probability that they are not consecutive ______.


6 boys and 6 girls sit in a row at random. The probability that all the girls sit together is ______.


The probability that the home team will win an upcoming football game is 0.77, the probability that it will tie the game is 0.08, and the probability that it will lose the game is ______.


A box contains 10 red marbles, 20 blue marbles and 30 green marbles. 5 marbles are drawn from the box, what is the probability that atleast one will be green?


If 4-digit numbers greater than 5,000 are randomly formed from the digits 0, 1, 3, 5, and 7, what is the probability of forming a number divisible by 5 when, the repetition of digits is not allowed?


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×